### v1

# ActionPose: Pretraining 3D Human Pose Estimation with the Dark Knowledge of Action

Abstract
2D-to-3D human pose lifting is an ill-posed problem due to depth ambiguity and occlusion. Existing methods relying on spatial and temporal consistency alone are insufficient to resolve these problems because they lack semantic information of the motions. To overcome this, we propose ActionPose, a framework that leverages action knowledge by aligning motion embeddings with text embeddings of fine-grained action labels. ActionPose operates in two stages: pretraining and fine-tuning. In the pretraining stage, the model learns to recognize actions and reconstruct 3D poses from masked and noisy 2D poses. During the fine-tuning stage, the model is further refined using real-world 3D human pose estimation datasets without action labels. Additionally, our framework incorporates masked body parts and masked time windows in motion modeling to mitigate the effects of ambiguous boundaries between actions in both temporal and spatial domains. Experiments demonstrate the effectiveness of ActionPose, achieving state-of-the-art performance in 3D pose estimation on public datasets, including Human3.6M and MPI-INF-3DHP. Specifically, ActionPose achieves an MPJPE of 36.7mm on Human3.6M with detected 2D poses as input and 15.5mm on MPI-INF-3DHP with ground-truth 2D poses as input.
Introduction
3D human pose estimation has been a significant research topic for better understanding human motion. It serves as a precursor for numerous downstream tasks such as human action recognition, human-computer interaction, and virtual reality. In monocular 3D human pose estimation, the input is derived from either 2D human poses or videos. However, converting 2D inputs to 3D outputs is an ill-posed problem due to depth ambiguity and occlusion, resulting in multiple possible 3D poses corresponding to a single 2D pose.
To address this challenge, recent work leverages spatial and temporal consistency. Various architectures, including TCN-based (Pavllo et al. 2019; Cheng et al. 2020), GCN-based (Cai et al. 2019; Ci et al. 2019; Wang et al. 2020), and transformer-based models (Zhang et al. 2022; Zhu et al. 2023; Peng, Zhou, and Mok 2024), have been designed to capture this information. And pretraining strategies involving random frame masking, joint masking, and noise addition on large-scale dataset have been proposed to enhance learning of spatial and temporal information (Zhu et al. 2023).
Figure 1: (Left) A side view of a person clapping, with overlapping hand joints. (Right) A front view of a person performing ‘Cloud Hands’ with the hands about to cross.
However, there exists many scenarios where spatial and temporal information alone is insufficient. For example, when a person claps their hands, the hands may overlap as they move toward each other from a side view, making it difficult to determine which hand is closer to the viewer, as shown in Figure 1. Such a depth ambiguity can be resolved with the knowledge of the clapping motion. Similarly, in the ‘Cloud Hands’ form of Tai Chi, one hand might always appear in front of the other from a front view despite that it actually moves back and forth relative to the other. Knowledge of the ‘Cloud Hands’ semantic context helps resolve this ambiguity. Thus, compared to spatial and temporal consistency, action information provides richer semantic context.
One intuitive way to integrate action clues into 3D human pose estimation is through multi-task learning: training 3D human pose estimation and human action recognition tasks simultaneously. However, very few datasets are available with both pose and detailed action descriptions. Most human pose estimation datasets only contain high-level action labels (e.g., Human3.6M (Ionescu et al. 2014) and MPI-INF-3DHP (Mehta et al. 2017)) and lack critical semantic information. To the best of our knowledge, to date, datasets with more descriptive action labels, such as BABEL (Punnakkal et al. 2021) and KIT (Krebs et al. 2021), suffer from significant class imbalance. For example, the verbs ‘walk’ and ‘transition’ occur frequently while the frequencies of other classes are significantly less in BABEL (Punnakkal et al. 2021). Moreover, textual action descriptions are inherently non-unique. Describing the timing of movements for different body parts is also challenging, especially during concurrent actions, resulting in ambiguous temporal and spatial boundaries.
This paper proposes a two-stage framework consisting of pretraining and fine-tuning to address the aforementioned challenges. During the pretraining stage, the model learns to recognize actions and reconstruct 3D poses from masked and noisy 2D poses. Rather than encoding action labels as one-hot vectors, as done in previous work (Luvizon, Picard, and Tabia 2018, 2021), we align the text embeddings of descriptive action labels with motion representations, leveraging the rich semantic information contained in actions. To tackle the issues of imbalanced data distribution in the pretraining dataset and the inherent uncertainty in human action recognition tasks, we replace the cross-entropy loss in infoNCE with a combination of KL-divergence and focal loss (Lin et al. 2017) to supervise multi-modal alignment. Additionally, inspired by MotionBERT (Zhu et al. 2023), we implement varying-size temporal window masking and body part-level masking during pretraining. This approach simulates scenarios where the spatial and temporal boundaries of an action are ambiguous and encourages the model to recognize the action even with partial information. By the end of the pretraining stage, the model is equipped to recognize actions. In the fine-tuning stage, the model is further refined using real-world 3D human pose estimation datasets without action descriptions.
In summary, the contributions of this work are three folds:
  • To the best of our knowledge, this is the first work to perform multi-modal pretraining that directly aligns the text embedding of descriptive action labels with motion sequences for a 3D human pose estimation task.
  • We propose a new pre-training and fine-tuning strategy for motion representation learning. By integrating multi-modal representation learning of motion with descriptive action labels and 3D human pose estimation, the model is trained to embed action cues into 3D pose estimation during the pretraining stage.
  • Experiments demonstrate the effectiveness of ActionPose, achieving state-of-the-art performance on public datasets. Specifically, ActionPose surpasses all existing methods, including those based on temporal information and diffusion-based approaches, on MPI-INF-3DHP. On Human3.6M, it significantly outperforms temporal information-based methods and achieves results comparable to diffusion-based methods with significantly lower inference overhead.
Related Work
3D Human Pose Estimation
3D human pose estimation can be categorized into two main approaches: the first involves estimating 3D human poses directly from images (Li and Chan 2015; Sun et al. 2018; Moon, Chang, and Lee 2019), and the second involves lifting 2D poses to 3D poses (Martinez et al. 2017; Liu et al. 2020; Zhang et al. 2022; Peng, Zhou, and Mok 2024). This work falls into the second category. Within this category, various architectures have been proposed to capture spatial and temporal information in motion dynamics, including TCN-based (Pavllo et al. 2019; Cheng et al. 2020), GCN-based (Cai et al. 2019; Ci et al. 2019; Wang et al. 2020), and transformer-based (Zhang et al. 2022; Zhu et al. 2023; Peng, Zhou, and Mok 2024) models.
Since action labels provide comprehensive summaries of motion, we adopt spatial and temporal transformers (Zhu et al. 2023) as our backbone for the pose encoder, which effectively learns and fuses motion representations across both temporal and spatial dimensions.
Skeleton-based Human Action Recognition
Another task related to understanding human motion is human action recognition. Unlike 3D human pose estimation, which focuses on precise pose reconstruction, human action recognition aims to comprehend human motion at a higher level. Despite the different objectives, both tasks require an understanding of the spatial relationships between the movements of different body joints and the temporal dynamics within motion. Consequently, recent works have designed various architectures to capture the spatio-temporal information of motion dynamics for human action recognition including transformer-based (Ahn et al. 2023; Zhu et al. 2023; Do and Kim 2024), GCN-based (Yan, Xiong, and Lin 2018; Chen et al. 2021; Chi et al. 2022; Zhou et al. 2024), LSTM-based (Zhu et al. 2016; Liu et al. 2017), and CNN-based (Zhang et al. 2019; Xu et al. 2022).
Figure 2: The overall architecture of ActionPose.
Multi-Task Learning
We use multi-task learning when the knowledge gained from one task can benefit other tasks (Zhang and Yang 2022). This is particularly relevant for human pose estimation and human action recognition. The work in (Luvizon, Picard, and Tabia 2018) was the first to highlight the interconnections between these tasks and attempted to estimate 2D/3D human poses alongside human action recognition. In their approach, both tasks share a common multi-task CNN, with human action recognition having an additional head to classify actions based on estimated poses and visual features. However, their method does not fully utilize the rich semantic information of actions, as the action classes are encoded as one-hot vectors. Similarly, (Luvizon, Picard, and Tabia 2021) aims to estimate 2D and 3D human poses and corresponding actions in real-time, but also encodes action classes as one-hot vectors.
Multi-Modal Representation Learning
Multi-modal representation learning methods, such as CLIP (Radford et al. 2021) and ALIGN (Jia et al. 2021), have proven highly effective for tasks like image captioning, text-image retrieval, and zero-shot image classification. These methods work by aligning text embeddings from a text encoder with feature embeddings from an image encoder. This alignment allows the image encoder to capture and utilize rich semantic information from images, significantly enhancing its performance on downstream tasks.
Pioneer works point out that similar techniques could be applied to human motion related tasks, such as human motion generation (Tevet et al. 2022), human action recognition (Xiang et al. 2023) and human pose estimation (Zheng et al. 2023). Among them, our work is most related to GAP (Xiang et al. 2023) and ActionPrompt (Zheng et al. 2023).
To enable each body part to recognize actions in human action recognition tasks, GAP (Xiang et al. 2023) aligns features of individual body parts with the corresponding text descriptions. The global action information is then derived from a combination of all parts of body, functioning like a model zoo where each part serves as an action predictor. In contrast, our proposed method is simple yet effective. We train each part of the body to classify actions by randomly masking them during training, similar to dropout techniques in machine learning. While some body parts are randomly omitted during training, all parts are utilized during inference.
Another work that aligns human motion and text embedding for human pose estimation is ActionPrompt (Zheng et al. 2023). Their approach consists of two components: an action-related text-prompt block and an action-specific pose-prompt block. The action-related text-prompt block helps the pose encoder recognize corresponding action labels from Human3.6M dataset (Ionescu et al. 2014), while the action-specific pose-prompt block refines poses based on the predicted action labels. However, because the action labels in Human3.6M dataset are coarsely defined and the learned text templates are shared across all actions, the semantic information within their text embeddings is not as rich as ours, which utilize action labels from the motion-with-language dataset (Punnakkal et al. 2021). Additionally, ActionPrompt (Zheng et al. 2023) requires storing pose-prompt embeddings for all actions, which is not memory efficient when the number of actions is large.
Methodology
Network Architecture
The overall objective of ActionPose is to enable 3D human pose estimation and text-motion alignment. The network consists of two encoders to extract features from text and pose data. Embedding pooling layers are then applied to generate global representations of the text and pose. These representations are used for text-motion alignment, and a regression head subsequently projects the pose features to estimate the 3D pose. The network architecture is illustrated in Figure 2.
Text and Pose Encoders
The ActionPose network consists of two parallel BERT-style models (Devlin et al. 2018) operating over pose and text domains. The text stream utilizes three layers of transformer blocks, following the architecture of the original BERT (Devlin et al. 2018). The pose stream employs five layers of Spatial-Temporal transformer blocks, following MotionBERT (Zhu et al. 2023).
Given a sequence of 2D pose skeletons $\mathbf{x}\in\mathbb{R}^{T\times J\times C_{in}}$, we first project them into a higher dimensional space through one MLP layer, resulting in features $\{p_{1},p_{2},...,p_{T}\}$ where $p_{i}\in\mathbb{R}^{J\times C_{f}}$, $C_{in}$ is the input dimension of 2D pose sequences, $C_{f}$ is the feature dimension of pose embeddings. We then concatenate a learnable pose class token to these features, forming the sequence $\{p_{0},p_{1},p_{2},...,p_{T}\}$ where $p_{0}$ is the class token <POSE>.
For the text input, we have $\{w_{0},w_{1},w_{2},...,w_{N}\}$, where $w_{0}$ is <CLS> and $w_{N}$ is <SEP>, namely tokens for classification and sentence separation. After processing the text inputs through the text encoder, our model generates encoded text features $\{h_{w_{0}}^{l_{3}},h_{w_{1}}^{l_{3}},h_{w_{2}}^{l_{3}},...,h_{w_{N}}^{l_{3}}\}$. Similarly, after processing the pose sequences through the pose encoder, our model produces encoded pose features $\{h_{p_{0}}^{l_{1}+l_{2}},h_{p_{1}}^{l_{1}+l_{2}},h_{p_{2}}^{l_{1}+l_{2}},...,h_{p_{T}}^{l_{1}+l_{2}}\}$.
Next, we introduce the structures to derive the global text representation and global pose representation, as follows.
Text Embedding and Pose Embedding Pooling Layers for Late Fusion
The purpose of these layers is to project the learned features of the pose class token <POSE> and the text class token <CLS> to the same feature space. These layers consist of multiple MLPs (Multi-Layer Perceptrons).
Instead of average pooling all pose features, we use the pose feature in the pose class token, as it represents a weighted summation of all pose features through attention. Since pose features include an additional joint dimension, we perform a learnable weighted summation over the joint dimension of the pose class token <POSE>.
We extract the global pose representation from the middle layer of the pose encoder $h_{{<\texttt{POSE}>}}^{l_{1}}$. Shallow layers are chosen because they capture more action-specific information, which can then be propagated to deeper layers for 3D human pose estimation. After this stage, we denote the text global representation as $h_{W}$ and the pose global representation as $h_{P}$ for simplicity. These global representations are used in the task of multi-modal alignment prediction.
Pose Estimation Regression Head
This structure is designed to project pose features $\{h_{p_{1}}^{l_{1}+l_{2}},h_{p_{2}}^{l_{1}+l_{2}},...,h_{p_{T}}^{l_{1}+l_{2}}\}$ to predict 3D pose sequence $\mathbf{X}\in\mathbb{R}^{T\times J\times 3}$.
Pretraining Tasks
We introduce two new pretraining strategies in addition to random joint and frame-level masked motion modeling: to address ambiguous temporal and spatial boundaries between actions and to improve the ability to classify actions using partial body poses.
Figure 3: The human body is partitioned into six parts for motion modeling with masked body parts.
Motion Modeling with Masked Body Parts
We partition the human body into six parts: upper body, lower body, head, arms, legs, and hips, as shown in Figure 3. One part is then randomly chosen to be masked.
In addition to enabling each body part to contribute to action classification, as discussed in Related WorkMulti-Modal Representation Learning and to mitigate the problem of ambiguous spatial boundaries between concurrent actions, this technique also helps the pose encoder learn the kinematics and anatomical details of the human body more comprehensively than joint-level masking alone, capturing the coordinated movements of multiple joints. When masked body part motion modeling is combined with multi-modal alignment prediction, the pose encoder further learns the spatial relationships of human motion within the context of corresponding actions. This masking technique also simulates real-world scenarios where parts of the human body may be occluded in a camera’s field of view.
Motion Modeling with Masked Time Windows
In addition to masked body parts, we propose another masking technique that masks the motion in the temporal domain with a varying window size between $[T_{1},T_{2}]$. The starting frame of this masked time window also varies and is randomly chosen.
The method can be interpreted as follows: masking different segments of motion in the time domain serves distinct purposes. Masking the initial part of the motion helps infer the cause of the action. Masking the middle segment allows us to model the transition between different motion states. Masking the latter part aids in predicting the motion’s intention and reconstructing human motion based on that predicted intention. Figure 4 provides a visualization of these interpretations.
Figure 4: Masking different segments of a human motion has distinct effects. Top: masking the initial part of the motion. Middle: masking the middle segment. Bottom: masking the latter part of the motion.
Multi-modal Alignment Prediction
The dual-stream encoders are optimized together by contrasting the global representation of text embeddings and pose embeddings at sample $j$:
$s_{p,w}^{j}=\frac{exp(h_{P}^{j}\cdot h_{W+}^{j}/\tau)}{\sum_{i=0}^{K}exp(h_{P}^{j}\cdot h_{W_{i}}^{j}/\tau)},$ (1)
where $h_{W+}^{j}$ is the corresponding text description of the action, $h_{W_{i}}^{j}$ for $i\in[0,1,...,K]$ includes one positive sample $h_{W+}^{j}$ and $K$ negative samples, and $\tau$ is the temperature.
Since motion-and-language datasets (Punnakkal et al. 2021; Krebs et al. 2021) are highly imbalanced and human action recognition tasks are inherently non-deterministic, namely, motion and action labels do not have a one-to-one correspondence, we replace the cross-entropy in the infoNCE loss with a combination of focal loss (Lin et al. 2017) and KL divergence. The modified loss function is as follows:
$\mathcal{L}_{con}=\sum_{j=1}^{M}(1-s_{p,w}^{j})^{\gamma}y_{j}log(\frac{y_{j}}{s_{p,w}^{j}}),$ (2)
where $M$ is the number of overall samples and $y_{j}$ is the ground-truth similarity score.
Objective Function
During pretraining, we adopt several masking strategies: joint-level and frame-level masking 50% of the time to capture spatial consistency among neighboring joints and temporal consistency across adjacent frames; random body part masking 25% of the time; and random time window masking 25% of the time. Additionally, we introduce noises into the 2D pose sequences.
Alongside the multi-modal alignment loss, we perform 3D pose reconstruction from the corrupted 2D poses. The discrepancy between the predicted 3D pose sequence $\mathbf{X}$ and the ground-truth 3D pose sequence $\hat{\mathbf{X}}$ is penalized using the following loss functions:
$\mathcal{L}_{3D}=\sum_{k=1}^{M}\sum_{t=1}^{T}\sum_{j=1}^{J}||\mathbf{\hat{X}_{t,j}^{k}}-\mathbf{X_{t,j}^{k}}||_{2},$ (3)
$\mathcal{L}_{v}=\sum_{k=1}^{M}\sum_{t=1}^{T}\sum_{j=1}^{J}||\mathbf{\hat{V}_{t,j}^{k}}-\mathbf{V_{t,j}^{k}}||_{2},$ (4)
Where $\mathbf{\hat{V}_{t}}=\mathbf{\hat{X}_{t}}-\mathbf{\hat{X}_{t-1}}$ and $\mathbf{{V_{t}}}=\mathbf{{X_{t}}}-\mathbf{{X_{t-1}}}$.
The final pretraining loss is computed by combining the multi-modal alignment loss and the reconstruction loss functions as follows:
$\mathcal{L}=\mathcal{L}_{con}+\lambda_{3D}\mathcal{L}_{3D}+\lambda_{v}\mathcal{L}_{v}.$ (5)
Fine-tuning
During pretraining, the parameters of both the pose encoder and text encoder are updated, but during fine-tuning, only the pose encoder is fine-tuned on the target dataset. Specifically, 2D poses are detected from videos and then lifted to 3D poses. The fine-tuning process is supervised by the reconstruction losses as specified in Equation 3 and Equation 4.
Method MPJPE P-MPJPE
P-STMO (Shan et al. 2022) 42.8 34.4
MixSTE (Zhang et al. 2022) 40.9 32.6
GLA-GCN (Yu et al. 2023) 44.4 34.8
STCFormer (Tang et al. 2023) 40.5 31.8
MotionBERT (Zhu et al. 2023$\diamond$ 37.5
DiffPose (Gong et al. 2023) †* 36.9
KTPFormer (Peng, Zhou, and Mok 2024) 40.1 31.9
KTPFormer (Peng, Zhou, and Mok 2024) †* 33.0 26.2
ActionPose (Ours) $\diamond$ 36.7 31.3
Table 1: Quantitative comparison of 3D human pose estimation on the Human3.6M dataset using detected 2D poses. Errors are reported as average MPJPE (mm) and P-MPJPE (mm). All methods use a temporal window of 243 frames. †indicates the use of temporal information, * represents diffusion-based methods, and $\diamond$ indicates pretraining-based methods. Results from other models are directly quoted from the respective papers.
Method PCK $\uparrow$ AUC $\uparrow$ MPJPE $\downarrow$
P-STMO (Shan et al. 2022) 97.9 75.8 32.2
GLA-GCN (Yu et al. 2023) 98.5 79.1 27.7
STCFormer (Tang et al. 2023) 98.7 83.9 23.1
DiffPose (Gong et al. 2023) †* 98.0 75.9 29.1
KTPFormer (Peng, Zhou, and Mok 2024) †* 98.9 85.9 16.7
MotionAGFormer-L (Mehraban, Adeli, and Taati 2024) 98.2 85.3 16.2
ActionPose (Ours) $\diamond$ 98.9 87.0 15.5
Table 2: Quantitative comparison of 3D human pose estimation on the MPI-INF-3DHP dataset. All methods use a temporal window of 81 frames. †indicates the use of temporal information, * represents diffusion-based methods, and $\diamond$ indicates pretraining-based methods. Results from other models are directly quoted from the respective papers.
Experiments
Pretraining
The BABEL (Punnakkal et al. 2021) dataset is a large dataset that includes language labels describing the actions performed in mocap sequences. It provides both sequence-level labels, which describe the overall action in the sequence, and frame-level labels, which detail fine-grained actions in every frame. Each frame in the sequence can have multiple corresponding fine-grained action labels. The motion sequences in the BABEL dataset are sourced from the large mocap dataset AMASS (Mahmood et al. 2019), which uses a common parameterization framework. Instead of obtaining videos and extracting 2D skeleton sequences from the mocap dataset, we project the 3D skeletons extracted from parametric models to 2D from both a side view and a front view, assuming an orthographic camera, as done in (Zhu et al. 2023). We align the body keypoint definitions with those of Human3.6M and convert the camera coordinates to pixel coordinates using the approach outlined in (Ci et al. 2022).
Each positive data pair for pretraining consists of a fine-grained frame-level action label and its corresponding motion sequence. For sequences labeled as “transition”, we concatenate the motion sequence before and after the “transition”, and use a template “transit from A to B” to construct new action labels. A and B are the corresponding action labels before and after the “transition”. Negative data pairs are generated by randomly associating action labels with unmatched motion sequences.
The overall pretraining framework consists of three layers of pretrained BERT (Devlin et al. 2018) as the text encoder and five layers of pretrained MotionBERT (Zhu et al. 2023) as the pose encoder. The network is trained on a single NVIDIA H100 GPU with a batch size of 16 and a sequence length of 243 frames and for 300 epochs.
Fine-tuning for 3D Human Pose Estimation
During fine-tuning for 3D human pose estimation, the text encoder is no longer required; only the pose encoder is fine-tuned on the target test dataset.
Datasets and Evaluation Metrics
We evaluated our methods on the public dataset Human3.6M (Ionescu et al. 2014) and MPI-INF-3DHP (Mehta et al. 2017).
Human3.6M (Ionescu et al. 2014) contains 3.6 million video frames of human motion performed by professional actors in an indoor environment. Following previous works (Zhu et al. 2023; Peng, Zhou, and Mok 2024), we use subjects 1, 5, 6, 7, and 8 for fine-tuning, and subjects 9 and 11 for testing. We report results using the mean per joint position error (MPJPE), which measures the average Euclidean distance between predicted and ground-truth joint positions, typically in millimeters. Additionally, we report the Procrustes-aligned MPJPE (P-MPJPE), where MPJPE is calculated after applying Procrustes alignment to the predicted and ground-truth positions.
MPI-INF-3DHP (Mehta et al. 2017) is another large-scale public dataset. This dataset contains recordings from 14 cameras capturing 8 actors performing 8 activities for the training set and 7 activities for evaluation. Following the setting of previous work (Shan et al. 2022; Mehraban, Adeli, and Taati 2024; Peng, Zhou, and Mok 2024), our evaluation metrics included the area under the curve (AUC), percentage of correct keypoints (PCK), and mean per-joint position error (MPJPE).
Performance Comparison on the Human3.6M Dataset
We evaluated ActionPose on the Human3.6M dataset (Ionescu et al. 2014) and reported MPJPE and P-MPJPE in millimeters. The 2D skeletons are extracted using Stacked Hourglass networks (Newell, Yang, and Deng 2016). The pose encoder is fine-tuned on the Human3.6M training set. As shown in Table 1, ActionPose significantly outperforms previous models based on temporal information and achieves results comparable to diffusion-based methods. Although the diffusion-based KTPFormer (Peng, Zhou, and Mok 2024) achieves better results on the Human3.6M dataset, diffusion-based methods are time-consuming during inference, as they require progressive refinement of the results.
Among all approaches, MotionBERT (Zhu et al. 2023) is the closest to us in its pretraining and fine-tuning framework. Table 1 shows that our model outperforms it by a large margin.
Performance Comparison on the MPI-INF-3DHP Dataset
To demonstrate the generalization ability of ActionPose, we evaluated its performance on the MPI-INF-3DHP dataset. Compared to Human3.6M, the test sets of MPI-INF-3DHP have more diverse motions and camera viewpoint variation.
We fine-tuned the pose encoder using ground-truth 2D poses as inputs, following the approach of previous work (Mehraban, Adeli, and Taati 2024; Peng, Zhou, and Mok 2024). As shown in Table 2, ActionPose achieves state-of-the-art results, with a PCK of 98.9%, an AUC of 87.0%, and an MPJPE of 15.5mm. Notably, these results surpass all existing state-of-the-art methods, including those based on temporal information and diffusion-based approaches.
Method MPI-INF-3DHP Human3.6M
Baseline 16.7 37.5
+ MMM 16.2 36.8
+ MAP 15.7 37.3
+ MMM + MAP 15.5 36.7
Table 3: Results of ablation study of each component of ActionPose on MPI-INF-3DHP and Human3.6M in MPJPE (mm). We refer to masked motion modeling as MMM and multi-modal alignment prediction as MAP.
Ablation Study
To verify the effectiveness of our proposed framework, we conducted ablation studies on the MPI-INF-3DHP dataset ($T$ = 81) using ground-truth 2D poses as inputs and on the Human3.6M dataset ($T$ = 243) using detected 2D poses. Table 3 presents the results, reported as MPJPE (mm), for each component added to our framework.
For the baseline, we used the pretrained MotionBERT model fine-tuned on the MPI-INF-3DHP and Human3.6M datasets, respectively. The pretrained MotionBERT weights were obtained from the official GitHub page. On MPI-INF-3DHP, pretraining with the proposed masked time windows and masked body parts motion modeling reduced the error by 0.5mm compared to the baseline. Incorporating random masking, random noise, and multi-modal alignment prediction further reduced the error by 1.0mm. Combining these strategies resulted in a total error reduction of 1.2mm from the baseline. Similarly, on Human3.6M, pretraining with masked time windows and masked body parts motion modeling reduced the error by 0.7mm, and adding random masking, noise, and multi-modal alignment reduced the error by an additional 0.2mm. The combination of these techniques led to a total error reduction of 0.8mm from the baseline.
Qualitative Results on Global Representation Learning
To verify ActionPose’s ability to learn rich semantic information about actions, we visualize the global representations of pose embeddings in Figure 5. Ideally, semantically similar actions should have closer embeddings, while different actions should be farther apart. For instance, ‘throw’ and ‘pick’ both emphasize upper body movements, and thus their representations are close. In constrast, ‘jump’, which emphasizes lower body movement, is distinctly separate from them. Similarly, ‘run’ and ‘kick’ are similar from leg movements and are positioned near each other, while ‘wave’ focused on the upper body, is more distant. Additionally, ‘bend’ and ‘raise’ are closely related, whereas ‘walk’ is farther from these actions.
As shown in Figure 5, ActionPose effectively groups semantically similar actions closer together while pushing semantically different actions farther apart. In contrast, MotionBERT’s global representation lacks a clear pattern, with significant overlap between both similar and distinct actions. Please note that there is some diversity within the same action due to variations in concurrent actions and different forms and styles of the actions. For example, “kicking while jumping” differs from “kicking while standing,” and the single action label with the verb ‘kick’ include motion sequences for both “kicking a soccer ball” or “kicking with the right leg back and up”.
(a) MotionBERT
(b) ActionPose
Figure 5: The t-SNE visualization of global representations of motion embeddings on the BABEL dataset: on the left (a) are embeddings obtained through random joint-level and frame-level masking, as proposed by MotionBERT (Zhu et al. 2023), with average pooling of all pose features in the sequence; on the right (b) are embeddings obtained through ActionPose.
Discussion
Learning sufficient action semantics during text embedding alignment with motion representation requires a large dataset. When the action classes in the pretraining dataset differ from those in the target dataset, the benefits of fusing action knowledge through pretraining may diminish. Despite this limitation, ActionPose can be extended to other human motion-related tasks, such as motion prediction and completion. For motion prediction, the model can use historical motion to predict intentions, thereby guiding future movements. Similarly, for motion completion, the model can learn transitions between starting and ending poses, constrained by the corresponding actions. The framework’s ability to learn rich semantic information about actions also has potential benefits for downstream tasks such as motion captioning and generation from text descriptions. These applications are reserved for future work.
Conclusion
In this work, we demonstrate how understanding human motion from an action perspective can enhance 3D human pose estimation. We propose a novel pretraining and fine-tuning strategy that embeds action cues during the pretraining stage. Experimental results confirm the effectiveness of the proposed masked motion modeling and multi-modal alignment tasks. Additionally, the ability to incorporate semantic information into global representation learning shows potential for benefiting downstream tasks such as human motion captioning, generation, and prediction.
Appendix
Additional Architecture Details
Figure 6: The architecture of pose embedding pooling layers (a) and text embedding pooling layers (b), where J represents the number of joints.
In this supplementary material, we provide a detailed illustration of the architecture for the Text Embedding and Pose Embedding Pooling Layers. As shown in Figure 6, all linear layers are activated by GELUs (Hendrycks and Gimpel 2023), followed by a layer normalization layer and/or a dropout layer. For the Pose Embedding Pooling layers, due to the extra joint dimension, we first reshape the input in the middle, then apply a weighted summation over the last dimension.
Experimental Details
Configuration
The network is implemented in PyTorch and trained on a single NVIDIA H100 GPU for 300 epochs using an AdamW optimizer, with a learning rate of 0.0005 and a batch size of 16.
Pretraining
For the pretraining datasets BABEL (Punnakkal et al. 2021) and AMASS (Mahmood et al. 2019), we first render the SMPL+H (Romero, Tzionas, and Black 2017) parametric model, then extract the 3D keypoints using a regression matrix (Bogo et al. 2016). The original AMASS dataset, which has a frequency of 120 Hz, is resampled to 30 Hz. In 50% of the random masking instances, 5% of the joints are masked at the joint level, and 15% of the frames are masked at the frame level. In 25% of the time-window masking instances, the size of the masking window is randomized between $T_{1}=30$ and $T_{2}=80$.
During pretraining, for the contrastive loss $\mathcal{L}_{con}$, the temperature is set to 0.1, and each pair of pose and text embeddings is compared against 16 negative samples.
Qualitative Results
Figure 7: Qualitative results of ActionPose compared to the baseline model MotionBERT. The left column shows the results from ActionPose, the middle column shows the results from MotionBERT for the same frames, and the right column displays the corresponding original 2D pose inputs.
Performance Comparison on the Human3.6M Dataset
In this section, we present qualitative results of our fine-tuned pose encoder on the Human3.6M dataset (Ionescu et al. 2014), comparing our results with the baseline model, MotionBERT (Zhu et al. 2023). As shown in Figure 7, ActionPose demonstrates robustness in cases of occluded 2D poses (first and third rows) and produces more accurate and natural-looking poses when 2D inputs are clear (second row).
Qualitative Results on the Pose Encoder’s Semantic Embedding
In this section, we present additional qualitative results to assess the semantic embedding capability of the fine-tuned pose encoder on the Human3.6M dataset (Ionescu et al. 2014). Specifically, we compare the similarity scores between pose embeddings and text embeddings of two selected action labels, followed by applying the SoftMax function to these scores. As shown in Figure 8, the pose encoder can distinguish not only between semantically similar actions (e.g., bend down and kneel) and semantically related actions (e.g., walk and stand), but also semantically inverse actions (e.g., sit down vs. get up from a chair, walk forward vs. walk backward).
Figure 8: Qualitative Results on the Pose Encoder’s Semantic Embedding. Each clip is manually selected from the Human3.6M dataset: (a) and (c) are from the activity ’Walking Dog’, while (b) is from ’Sitting Down’.